The reaction A + B → C + D is second order in A and zero order in B. The value of k is 0.012 M-1min-1. What is the rate of this reaction when [A] is 0.125 M and [B] = 0.435 M?

A. 5 x 10-4

B. 3.4 x 10-3

C. 1.9 x 10-4

D. 1.3
Answer #1

At B C to Given; Rate = k [A]”[6]° = K[A]? [A] = 0.125 m; K= 0·012 & Rate = 0.012 + (0.125)² | Rate = 109x104 In mint